In a challenging chess game, the probability that Mr.B will win is 0.67. Assume that the chess competition follows a binomial distribution and 7 separate matches will be played. 

Determine: 


(a) will win at most 3 matches 

(b) will lose exactly 5 matches 

(c) will win every match


1
Expert's answer
2022-02-18T05:35:17-0500

Let "X=" the number of matches won by Mr.B: "X\sim Bin(n, p)."

Given "n=7, p=0.67, q=1-p=1-0.67=0.33"

(a)


"P(X\le 3)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=3)"

"=\dbinom{7}{0}(0.67)^0(0.33)^{7-0}+\dbinom{7}{1}(0.67)^1(0.33)^{7-1}"

"+\dbinom{7}{2}(0.67)^2(0.33)^{7-2}+\dbinom{7}{3}(0.67)^3(0.33)^{7-3}"

"=0.00042618443+0.00605698477"

"+0.03689254363+0.1248384052"

"=0.168214"

(b)


"7-5=2"

"P(X=2)=\dbinom{7}{2}(0.67)^2(0.33)^{7-2}"

"=0.036893"

(c)


"P(X=7)=\dbinom{7}{7}(0.67)^7(0.33)^{7-7}"

"=0.060607"